A 400 mH coil of negligible resistance is connected to an AC circuit in which an effective current of 6 mA is flowing. Find out the voltage across the coil if the frequency is 1000 Hz


1
Expert's answer
2021-03-02T18:06:20-0500

The reactance of the coil is given as follows:


"X_L = \omega L"

where "L = 400mH = 0.4H" is the inductance of the coil, and


"\omega = 2\pi f"

is the angular frequency of the source ("f = 1000Hz").

According to the Ohm's law, the effective voltage across the coil is:


"U = IX_L = 2\pi fLI"

where "I = 6mA = 0.006A" is the effective current through the coil. Thus, obtain:


"U = 2\pi\cdot 1000\cdot 0.4\cdot 0.006 \approx 15.08V"

Answer. 15.08 V.